In Newton’s rings experiment using light of two wavelengths , it is observed that 10 th order bright ring corresponding to wavelength 1 coincides with 15 th order dark ring corresponding to wavelength λ₂ when viewed by reflected light. Find the ratio of λ1 and λ₂


1
Expert's answer
2022-03-24T16:37:12-0400

"r_1=\sqrt{\frac{m_1R\lambda_1}{n}},"

"r_2=\sqrt{\frac{m_2R\lambda_2}{2n}},"

"r_1=r_2,"

"\frac{\lambda_1}{\lambda_2}=\frac{m_2}{m_1}=1.5."